# Marco-Voice Technical Report

### v1

@@ Abstract
This paper presents a multifunctional speech synthesis system that integrates voice cloning and emotion control speech synthesis within a unified framework. The goal of this work is to address longstanding challenges in achieving highly expressive, controllable, and natural speech generation that faithfully preserves speaker identity across diverse linguistic and emotional contexts. Our approach introduces an effective speaker-emotion disentanglement mechanism with in-batch contrastive learning, enabling independent manipulation of speaker identity and eemotional style, as well as rotational emotional embedding integration method for smooth emotion control. To support comprehensive training and evaluation, we construct CSEMOTIONS, a high-quality emotional speech dataset containing 10 hours of Mandarin speech from six professional speakers across seven emotional categories. Extensive experiments demonstrate that our system, Marco-Voice, achieves substantial improvements in both objective and subjective metrics. Comprehensive evaluations and analysis were conducted, results show that Marco-Voice delivers competitive performance in terms of speech clarity and emotional richness, representing a substantial advance in the field of expressive neural speech synthesis.
@@ 1 Introduction
Despite these achievements, truly human-like speech synthesis remains an open challenge [Li et al., 2024a]. In natural communication, human speech is characterized by a rich interplay of speaker identity (timbre), prosodic style (intonation, rhythm, emphasis), and nuanced emotional expression [Tan et al., 2021, Zhang et al., 2023, Barakat et al., 2024]. Replicating this diversity and flexibility in synthetic speech requires effective modeling and disentanglement of these factors [Wang et al., 2024, Meng et al., 2025]. The motivation for this work stems from three persistent challenges in the field: 1) Entanglement of Timbre and Speaking Style: Many TTS models intertwine speaker-specific timbre with prosodic style [Chen et al., 2024c], making it difficult to independently control voice identity and manner of speaking. This limitation restricts the personalization and expressiveness of synthesized voices, particularly in applications that require voice cloning or style transfer. 2) Balancing Prosody and Emotion Consistency: Achieving both natural prosody and consistent, expressive emotional content is difficult [Wu et al., 2019, Li et al., 2022]. Systems often excel at one aspect at the expense of the other, resulting in speech that sounds either monotonic or emotionally incongruent [Li et al., 2024c]. 3) Limitations of Conventional Emotion Modeling: Most existing TTS systems represent emotions using discrete categories (e.g., happy, sad, angry), which fails to capture the continuous and multidimensional nature of real-world emotional expression. Moreover, these methods often struggle to maintain high speaker similarity when synthesizing emotional speech, especially in voice cloning scenarios [Li et al., 2021, Kansizoglou et al., 2022]. 4) Limited Availability of High-Quality Emotional Speech Data: Existing emotional speech datasets often suffer from limited speaker diversity, inconsistent recording conditions, or insufficient emotional coverage, particularly for non-English languages [Tits et al., 2020, Zhou et al., 2022, Ma et al., 2024], which constrains the development and evaluation of emotional TTS systems.
@@ 2.3 Speaker-Emotion Disentanglement
[…]
@@
[…]
@@ 2.5 Conditional Flow Matching Module
$\begin{gathered}Q=W_{q}(\mathbf{e})\\ K=W_{k}(\mathbf{h}_{LM})\\ V=W_{v}(\mathbf{h}_{LM})\\ \end{gathered}$ (6)
@@ 3.1 Datasets
  • CSEMOTIONS (Chinese Speech Emotions): To supplement existing public resources, we constructed CSEMOTIONS, a proprietary emotional speech dataset. This corpus includes about 10 hours of Mandarin speech from six professional native speakers (three male, three female), all with extensive voice acting experience. CSEMOTIONS covers seven distinct emotional categories, with each speaker reading a curated set of 100 Chinese and English prompts. All recordings were made in professional studios to ensure high fidelity and expressive consistency.
@@ 4.1 Voice Cloning Evaluation
As shown in Table 1, Marco-Voice consistently outperforms existing voice cloning systems across all evaluated dimensions. Notably, our system achieves the highest speaker similarity score (0.8275), demonstrating its effectiveness in preserving speaker identity. Improvements in speech clarity, rhythm, and overall satisfaction further highlight the advantages of our timbre-style disentanglement approach.
@@ 4.4 Analysis Studies
Table 4: Objective evaluation of speech recognition and synthesis quality on the LibriTTS dataset. Metrics include word error rate (WER), speaker similarity (SS) using SpeechBrain and ERes2Net, error counts (Del & Ins, Sub), and DNS-MOS for perceptual quality. Lower WER and error counts, and higher SS and DNS-MOS indicate better performance.

### v3

@@ Abstract
This paper presents a multifunctional speech synthesis system that integrates voice cloning and emotion control speech synthesis within a unified framework. The goal of this work is to address longstanding challenges in achieving highly expressive, controllable, and natural speech generation that faithfully preserves speaker identity across diverse linguistic and emotional contexts. Our approach introduces an effective speaker-emotion disentanglement mechanism with in-batch contrastive learning, enabling independent manipulation of speaker identity and emotional style, as well as rotational emotional embedding integration method for smooth emotion control. To support comprehensive training and evaluation, we construct CSEMOTIONS, a high-quality emotional speech dataset containing 10 hours of Mandarin speech from six professional speakers across seven emotional categories. Extensive experiments demonstrate that our system, Marco-Voice, achieves substantial improvements in both objective and subjective metrics. Comprehensive evaluations and analysis were conducted, results show that Marco-Voice delivers competitive performance in terms of speech clarity and emotional richness, representing a substantial advance in the field of expressive neural speech synthesis. Our code and dataset are publicly available at https://github.com/AIDC-AI/Marco-Voice and https://huggingface.co/datasets/AIDC-AI/CSEMOTIONS respectively.
@@ 1 Introduction
Despite these achievements, truly human-like speech synthesis remains an open challenge [Li et al., 2024a]. In natural communication, human speech is characterized by a rich interplay of speaker identity, prosodic style (intonation, rhythm, emphasis), and nuanced emotional expression [Tan et al., 2021, Zhang et al., 2023, Barakat et al., 2024]. Replicating this diversity and flexibility in synthetic speech requires effective modeling and disentanglement of these factors [Wang et al., 2024, Meng et al., 2025]. The motivation for this work stems from three persistent challenges in the field: 1) Entanglement of Emotion and Speaking Style: Many TTS models intertwine speaker-specific emotion with prosodic style [Chen et al., 2024c], making it difficult to independently control voice identity and manner of speaking. This limitation restricts the personalization and expressiveness of synthesized voices, particularly in applications that require voice cloning or style transfer. 2) Balancing Prosody and Emotion Consistency: Achieving both natural prosody and consistent, expressive emotional content is difficult [Wu et al., 2019, Li et al., 2022]. Systems often excel at one aspect at the expense of the other, resulting in speech that sounds either monotonic or emotionally incongruent [Li et al., 2024c]. 3) Limitations of Conventional Emotion Modeling: Most existing TTS systems represent emotions using discrete categories (e.g., happy, sad, angry), which fails to capture the continuous and multidimensional nature of real-world emotional expression. Moreover, these methods often struggle to maintain high speaker similarity when synthesizing emotional speech, especially in voice cloning scenarios [Li et al., 2021, Kansizoglou et al., 2022]. 4) Limited Availability of High-Quality Emotional Speech Data: Existing emotional speech datasets often suffer from limited speaker diversity, inconsistent recording conditions, or insufficient emotional coverage, particularly for non-English languages [Tits et al., 2020, Zhou et al., 2022, Ma et al., 2024], which constrains the development and evaluation of emotional TTS systems.
@@ 2.3 Speaker-Emotion Disentanglement
$\text{Dot-Product Matrix:}\quad D=ES^{T}$
$\text{Norms:}\quad n_{E}=\|E\|,\quad n_{S}=\|S\|$
$\text{Normalized Matrix:}\quad\hat{D}=D/(n_{E}n_{S}^{T})$
@@
where $B$ is batch size, $S$ and $E$ consists of a batch of $\mathbf{s}$ and $\mathbf{e}$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $\text{cos\_sim}(E,S)$ computes the cosine similarity between each emotion embedding and each speaker embedding. During training, if batchwise pairwise computation is enabled, we average the absolute dot product between each embedding and all opposing emotion embeddings in the batch, except self-pairs; otherwise, we use the orthogonality loss as above. This constraint forces speaker and emotion embeddings to be perpendicular in the feature space, enabling independent control over voice identity and emotional expression.
@@ 2.5 Conditional Flow Matching Module
[…]
@@ 3.1 Datasets
  • CSEMOTIONS (Chinese Speech Emotions): To supplement existing public resources, we constructed CSEMOTIONS, a proprietary emotional speech dataset. This corpus includes about 10 hours of Mandarin speech from ten professional native speakers (five male, five female), all with extensive voice acting experience. CSEMOTIONS covers seven distinct emotional categories, with each speaker reading a curated set of 100 Chinese and English prompts. All recordings were made in professional studios to ensure high fidelity and expressive consistency.
@@ 4.1 Voice Cloning Evaluation
As shown in Table 1, Marco-Voice consistently outperforms existing voice cloning systems across all evaluated dimensions. Notably, our system achieves the highest speaker similarity score (0.8275), demonstrating its effectiveness in preserving speaker identity. Improvements in speech clarity, rhythm, and overall satisfaction further highlight the advantages of our speaker-style disentanglement approach.
@@ 4.4 Analysis Studies
Table 4: Objective evaluation of speech recognition and synthesis quality on the LibriTTS dataset. Metrics include word error rate (WER), speaker similarity (SS) using SpeechBrain and ERes2Net, error counts (Del & Ins, Sub), and DNS-MOS for perceptual quality. Lower WER and error counts, and higher $S$ and DNS-MOS indicate better performance.